OpenStudy (anonymous):

1. A dentist made the scatterplot below to show the number of cavities her patients had as it relates to the number of times they flossed their teeth each week. Which of the following best describes the correlation for the data? (1 point) Positive correlation Nonlinear correlation Negative correlation No correlation 2. What is the vertex of the graph of the following quadratic function: f(x) = x2 + 6x + 10 (1 point) (3, -1) (-3, -1) (-3, 1) (3, 1) 3. Which of the following graphs best represents a system of equations that has no solution? (1 point)

OpenStudy (anonymous):

3. Which of the following graphs best represents a system of equations that has no solution? (1 point) 4. The graph below shows the relationship between the number of dollars a worker earns and the number of hours worked. What does the slope of the graph represent? (1 point) The number of hours of work it takes to earn $320. The amount of money earned per hour. The amount earned for 40 hours of work. The number of hours worked per dollar earned. 5. The perimeter of a rectangle is 42 centimeters. The length of the rectangle can be represented by (x + 4), and its width can be represented by (2x - 7). What are the dimensions of this rectangle in centimeters? (1 point) Length = 10 and Width = 11 Length = 8 and Width = 13 Length = 6 and Width = 15 Length = 12 and Width = 9 6. A graph is shown below. Which of the following equations are represented by the graph? I. y = -x - 2 II. 2x - 3y = 6 III. y = (x - 2)(x - 3 ) IV. y - 2 = 2(x - 6) (1 point) II and IV I and III II and III I and IV 7. Which table shows the same relationship as y = -x2 + 3x ? (1 point) 8. Which graph shows a function with a domain of all real numbers greater than 7? (1 point) 9. What is the solution set for the quadratic equation x2 - 16 = 0 ? (1 point) {4} {-4, 4} {256} {-256, 256} 10. A painter charges $35 per hour for labor plus $40 for a ladder rental when he paints a house. The customer provides the paint. The total charge to paint a customer’s house was $950. How many hours did the painter spend painting this house? (1 point) h 28 h 23 h Not here 11. The slope and y-intercept of the graph of f were changed to make the graph of g, as shown below. Which statement describes the changes that were made to the graph of f to make the graph of g? (1 point) The slope was multiplied by 2, and the y-intercept was decreased by 5 to make the graph of g. The slope was multiplied by , and the y-intercept was increased by 5 to make the graph of g. The slope was multiplied by -2, and the y-intercept was decreased by 5 to make the graph of g. The slope was multiplied by , and the y -intercept was increased by 5 to make the graph of g . 12. The population of a town is currently 9,000 . The function p = 9,000 + 8t2 can be used to estimate p, the population of the town t years from now. Based on this function, which statement is true? (1 point) The population of the town is increasing at a constant rate. The population of the town will reach 10,000 between 11 and 12 years from now. The population of the town will increase by 256 people two years from now. The population of the town will increase and then decrease. 13. The late fee for overdue books at a library is $0.25 per day per book, with a maximum late fee of $5.00 per book. Which graph models the total late fee for 3 books that were checked out on the same day and are overdue? (1 point) 14. Which inequality is equivalent to 7x - 2y > 8? (1 point) y > x + 8 y > x + y < x - 4 y < x - 15. An airplane’s altitude in feet during its descent for landing can be found using the function f(x) = -300x + 30,000, where x represents the horizontal distance in miles from where the plane begins its descent. After new government regulations become law, the airplane’s descent will be modeled by the function g(x) = -300x + 30,500. Which statement describes this change? (1 point) The airplane starts its descent from an altitude 500 feet higher. The airplane starts its descent from an altitude 500 feet lower. The airplane descends 500 feet per horizontal mile faster. The airplane descends 500 feet per horizontal mile slower. 16. The graph of the quadratic function h passes through the points (-4, 32), (3, 4), (5, 14), and (7, 32). Which of the following shows the same relationship as h? (1 point) h(x) = x2 + 3x + 4 17. The function y = x2 + x - 6 is graphed below. What are the values of x when x2 + x - 6 = -4? (1 point) x = -4 and x = 6 x = -2 and x = 1 x = -3 and x = 2 x = -5 and x = -6 18. Which coordinate pair is in the solution set for y < 1 - 6x? (1 point) (1, 0) (1, -1) (0, 1) (-1, 1) 19. The graph of line p represents y = x - 1. If the slope of line p is multiplied by -10 to create line r, which statement about the graphs of the two lines is true? (1 point) Line r intersects line p. Line r is parallel to line p. Line r is 10 units above line p. Line r is 10 units below line p. 20. A teacher collected data on 20 students for two different quizzes. The scatterplot below shows the relationship between the number of points scored on Quiz 1 and the number of points scored on Quiz 2. Which statement describes the data? (1 point) The number of points scored on Quiz 2 was less than the number of points scored on Quiz 1 for any student who scored at least 50 points on Quiz 1.

OpenStudy (anonymous):

The number of points scored on Quiz 2 was greater than the number of points scored on Quiz 1 for any student who scored at least 50 points on Quiz 1. The number of points scored on Quiz 2 was less than the number of points scored on Quiz 1 for any student who scored 50 or fewer points on Quiz 1.

OpenStudy (anonymous):

21. A high school band held a bake sale. The number of cupcakes sold was four more than twice the number of cookies sold. The band sold a total of 52 cupcakes and cookies. How many cupcakes were sold? (1 point) 28 16 36 24 22. Two points on the graph of a quadratic function are shown on the grid below. What is the equation for the axis of symmetry of the graph of this function? (1 point) x = -3 y = -3 x = -5 y = -5 23. The graph below shows the water level in a tank being drained at a constant rate. If the rate at which the tank is drained is changed to 3 inches per hour and the initial water level stays the same, how would the time it takes to empty the tank be affected? (1 point) It would take 4 fewer hours. It would take 1.5 more hours. It would take 2 fewer hours. It would take 2 more hours. 24. A weightlifter is adding plates of equal weight to a bar. The table below shows the total weight, including the bar, that he will lift depending on the total number of plates on the bar. Based on this information, which statement is true? (1 point) The bar weighs 35 lb without any plates. The bar weighs 70 lb without any plates. The bar weighs 45 lb without any plates. The bar weighs 25 lb without any plates. 25. What is the range of the function shown below? (1 point) {-7, -2, 0.5} {-9, -4, -1} {-9, -7, -4, -2, -1, 0.5} {-1} 26. Which set of ordered pairs contains only points that are on the graph of the function y = 12 - 3x? (1 point) {(-3, -27), (0, 0), (6, 54)} {(-18, 10), (-6, 6), (18, -2)} {(-5, 27)}, (-1, 15), (8, -12)} {(-7, -9), (-4, 0), (2,18)} 27. Which inequality can be represented by the graph below? (1 point) y ≥ x + b x - y ≥ -b x + y ≤ b -y ≤ x + b 28. If y = x -2 , what is the value of x when y = -9? (1 point) 29. Which expression is equivalent to -6x2 - 11x - 4? (1 point) (3x + 7)(3x - 3) (-3x + 4)(2x - 1) (3x - 7)(3x + 3) (-3x - 4)(2x + 1) (3x - 7)(3x + 3) (-3x - 4)(2x + 1)

Directrix (directrix):

On #1, please post the scatterplot --> "made the scatterplot below" Also, please post only one question per thread. I am not saying that you are, but when a list of 29 questions is posted, it gives the appearance that you are asking Answerers to do your homework for you. We are not allowed to give only answers. Our mission is to work with you to arrive at a correct answer. @lolanswers

OpenStudy (anonymous):

I just wanna check my work plz and explanation plz help me out

OpenStudy (anonymous):

From connection academy any body help I will give medals